# 1.27in1.27in1in1in

### v1

@@ 1. Introduction
Based on our description, one can implement the Mori fan on a computer. This was the original motivation for this work, but in the course of the project, it became clear that there were theoretical questions to be answered that are equally interesting. Apart from the main theorem, this includes e.g. a better understanding of interior special points as in Section 3 or an explicit description of the cone of curves in terms of the curve structure, see Proposition 5.3. As a follow-up project, we have implemented combinatorial K3 surfaces in SageMath and will use it in forthcoming work to further study the Mori fan and related constructions.
@@ Acknowledgements
Klaus Hulek was partially supported by DFG grant Hu 337/7-2. Christian Lehn was supported by the DFG through the research grants Le 3093/3-2 and Le 3093/5-1.
@@ A $d$-semistable K3 surface of type III is a normal crossing surface $Y=\bigcup Y_{i}$ such that
(2.1) $0\xrightarrow{\ \ }\operatorname{Pic}(Y)\xrightarrow{\ \ }\bigoplus_{i}H^{2}(Y_{i}^{\nu},\mathbb{Z})\xrightarrow{\ \ }\bigoplus_{C}H^{2}(C,\mathbb{Z})$
where $C$ runs through all $1$-dimensional components of preimages of the double curves, see equation (2.3) on p.9 of [HL22]. The following definition is central to this article.
@@
[…]
@@
[…]
@@
A combinatorial K3 surface $Y$ is said to be in $(-1)$-form if the following holds. For any integral curve $D\subset Y^{\nu}$ in the preimage of the double locus, we have $D^{2}=-1$ if $D$ is smooth and $D^{2}=1$ if $D$ is singular (and hence a nodal rational curve), cf. the corresponding notion for models in [HL22, Definition 2.6].
@@
Taking the dual intersection complex defines a bijection between the set isomorphism classes of combinatorial K3 surfaces in $(-1)$-form and triangulations of the sphere ${\mathbb{S}}^{2}$ such that no vertex has valency greater than $6$. In particular, a combinatorial K3 surface in $(-1)$-form is uniquely determined by its dual intersection complex.∎
@@
Let $G$ be a triangulation of the sphere ${\mathbb{S}}^{2}$. We denote by $Y_{G}$ the unique combinatorial K3 surface in $(-1)$-form with dual intersection complex $G$.
@@
The classification of combinatorial K3 surfaces in $(-1)$-form through special triangulations of the sphere from Proposition 2.7 is a first instance of the combinatorial nature of combinatorial K3 surfaces. But not only those in $(-1)$-form are determined by combinatorial objects (at least in degree $2$, even though this is likely to hold true in higher degree as well): a general combinatorial K3 surface is determined by its curve structure, a certain graph that captures the configuration of curves on the surface, see Definition 2.19. This fully justifies the terminology combinatorial K3 surface.
@@ 2.2. Type I flops
Assume we have a smooth curve $C_{i}\subset Y_{i}$ in one of the components of $Y$, whose preimage is a $(-1)$ curve on the normalization $Y^{\nu}_{i}$, and which intersects the double locus in one point $P_{i}\in D_{ij}$, which is not a triple point, and where the intersection is transversal. Here $i=j$ is allowed, in which case the assumption is that in the normalization, $C_{i}^{\nu}$ meets one of the components of the preimage $D_{ii}^{1}\cup D_{ii}^{2}$, say $D_{ii}^{1}$. We then perform the following construction.
If $i\neq j$, we set $Y^{\prime\nu}_{k}:=Y^{\nu}_{k}$ for $k\neq i,j$. We have also natural identifications [^2] Let us point out that these identifications will preserve so-called interior special points, see Definition 3.2. This is however not relevant for the construction of the flop in the type I case. of the components $D_{ki},D_{kj}$ and $D^{\prime}_{ki},D^{\prime}_{kj}$ of the components of the anticanonical cycle. We define $Y^{\prime\nu}_{i}$ as the blow-down of $Y_{i}^{\nu}$ along the preimage of the $(-1)$-curve $C_{i}$. We further define $Y^{\prime\nu}_{j}$ as the blow-up of $Y_{j}$ in the preimage [^3] This will be called the interior special point on $D_{ji}$ in the next section. of the point where $C_{i}$ meets $Y_{j}$.
@@
2.3. Combinatorial K3 surfaces of degree $2$
In this article, we will mainly be concerned with combinatorial K3 surfaces of degree $22d=$ 2. Such surfaces have $d+1=3$ irreducible components, and the dual intersection complex is one of the two triangulations of ${\mathbb{S}}^{2}$ with three vertices, see Figure 2. As indicated in the figure, we will denote these triangulations by $P$ respectively $T$. Depending on its dual intersection complex, a combinatorial K3 surface of degree $2$ is thus said to be of class $P$ resp. $T$.
[…]
@@ 2.3. Combinatorial K3 surfaces of degree $2$
By Proposition 2.7, there are two combinatorial K3 surfaces of degree $2$ in $(-1)$-form, one each of class $P$ resp. $T$. Following [HL22, Section 2.2], we call them $Y_{P}$ and $Y_{T}$. Let us recall their construction and some basic facts about their geometry.
@@
By [Laz08] (see also [HL22, Construction 2.25] for more details and precise references), there are weak del Pezzo surfaces ${\mathfrak{Y}}_{i}$, $i=1,\ldots,6$ such that any connected component of the normalization of any $Y\in{\operatorname{CK3}^{\scriptscriptstyle(-1)}}$ is isomorphic to one of those. More precisely, the ${\mathfrak{Y}}_{i}$ come with an anticanonical cycle ${\mathfrak{D}}_{i}$ having $i$ irreducible components and the components of the normalization of $Y$ are isomorphic to $({\mathfrak{Y}}_{i},{\mathfrak{D}}_{i})$ as log Calabi–Yau surfaces, see Remark 2.2. While the details of their construction can be found in the above references, we are mainly interested in their combinatorial description via curve structures. The latter will be defined in general in Section 2.4, but for the surfaces ${\mathfrak{Y}}_{i}$ it is straightforward: we define the set
$C({\mathfrak{Y}}_{i}):=\{C\subset{\mathfrak{Y}}_{i}\mid C\text{ is an integralcurve with }C^{2}<0,C\not\subset{\mathfrak{D}}_{i}\}.$
@@
For us, only the surfaces ${\mathfrak{Y}}_{1}$, ${\mathfrak{Y}}_{2}$, and ${\mathfrak{Y}}_{4}$ will play a role. They show up in the construction of the models in $(-1)$-form, $Y_{P}$ and $Y_{T}$, as we will explain next.
$-1$
@@
11-1- $-1$
Figure 4. The augmented curve structure of the sruface ${\mathfrak{Y}}_{2}$.
11-1- 111-1- 111-1- $-1$
@@
The surface $Y_{P}$ has three identical irreducible components $Y_{1}=Y_{2}=Y_{3}={\mathfrak{Y}}_{2}$ whose augmented curve structure is shown in Figure 4. The black vertices represent the components of the anticanonical cycle and the white vertices have self-intersection $(-2)$ unless otherwise indicated. Call $D_{ij}$, $D_{ik}$ with $\{i,j,k\}=\{1,2,3\}$ the boundary components of $Y_{i}$. Then $Y_{P}$ is obtained by gluing the $Y_{i}$ cyclically along the isomorphisms $D_{ij}{\ \cong\ }D_{ji}$ for all $i,j$.
@@
The surface $Y_{T}$ has irreducible components $Y_{1}=Y_{2}={\mathfrak{Y}}_{1}$ and $Y_{3}$ a nonnormal surface with normalization ${\mathfrak{Y}}_{4}$. We call $Y_{3}$ the special component and $Y_{1}$, $Y_{2}$ the nonspecial components. The curve structures of $Y_{1}=Y_{2}={\mathfrak{Y}}_{1}$ and $(Y_{3})^{\nu}={\mathfrak{Y}}_{4}$ are shown in Figure 3 respectively Figure 5. To obtain the surface $Y_{T}$, one glues the components as follows. First, one glues two opposite boundary components to one another, e.g. those represented by the black vertices in the bottom row of Figure 5, such that the intersection points with the remaining boundary components are matched and also the intersection points with the curves corresponding to unique white vertices adjacent to the glued boundary components are matched. This specifies the gluing uniquely. After this first gluing, the remaining boundary components, i.e. those represented by the black vertices in the top row of Figure 5, have become nodal rational curves. These are then glued to the boundary curves of the nonspecial components, again matching singular points and intersection points with neighboring curves in the curve structure.

### v2

# Combinatorial K3 surfaces and the Mori fan of the Dolgachev–Nikulin–Voisin family in degree 2

@@ 1. Introduction
Based on our description, one can implement the Mori fan on a computer. This was the original motivation for this work, but in the course of the project, it became clear that there were theoretical questions to be answered that are equally interesting. Apart from the main theorem, this includes e.g. a better understanding of interior special points as in Section 3 or an explicit description of the cone of curves in terms of the curve structure, see Proposition 5.3. As a follow-up project, we have implemented combinatorial K3 surfaces in SageMath [Sage24] and will use it in forthcoming work to further study the Mori fan and related constructions.
@@ Acknowledgements
Klaus Hulek was partially supported by DFG grant Hu 337/7-2. Christian Lehn was supported by the DFG through the research grants Le 3093/3-2 and Le 3093/5-1. We thank the referee for their thorough reading and valuable comments.
@@ A $d$-semistable K3 surface of type III is a normal crossing surface $Y=\bigcup Y_{i}$ such that
(2.1) $0\xrightarrow{\ \ }\operatorname{Pic}(Y)\xrightarrow{\ \ }\bigoplus_{i}H^{2}(Y_{i}^{\nu},\mathbb{Z})\xrightarrow{\ \ }\bigoplus_{C}H^{2}(C^{\nu},\mathbb{Z})$
where $C$ runs through all double curves, see equation (2.3) on p.9 of [HL22]. Note that the curve $C$ might get normalized in $Y_{i}^{\nu}$, so it is more natural to write $C^{\nu}$ in the last term, even if it does not change anything on the level of $H^{2}$. The following definition is central to this article.
@@
Note that the notion of maximality from the previous definition coincides with the one from [HL22, Definition 2.15]. This is shown in [HL22, Lemma 2.10], taking into account equation (2.4) in op. cit. In particular, the central fiber of a type III degeneration is a maximal $d$-semistable K3 surface if and only if the degeneration itself is maximal.
@@
    Equivalently, we can describe a combinatorial K3 surface as follows. Instead of a surface $Y$ as above, we describe it as a semisimplicial object ${Z_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z_{0}}$ of length $2$. Here, $Z_{0}=Y^{\nu}$ is the normalization and $Z_{1}$ is the disjoint union over all double curves. One can think of this as a gluing datum, i.e. $Y$ as the pushout of this diagram. Note that pushouts always exist for finite morphisms by [Fer03, Théorème 5.4].
@@
$\operatorname{CK3}^{\scriptscriptstyle(-1)}$}:
The subclass of $\operatorname{CK3}_{2d}$} consisting of surfaces in (1)(-1)-form.
We write $\operatorname{CK3}_{2d}$} and $\operatorname{CK3}_{2d}^{\scriptscriptstyle(-1)}$} for the subclasses of surfaces of degree $2d$.
@@
A combinatorial K3 surface $Y$ is said to be in (1)(-1)-form if the following holds. For any integral curve $D\subset Y^{\nu}$ in the preimage of the double locus, we have $D^{2}=-1$ if $D$ is smooth and $D^{2}=1$ if $D$ is singular (and hence a nodal rational curve), cf. the corresponding notion for models in [HL22, Definition 2.6].
@@
Taking the dual intersection complex defines a bijection between the set of isomorphism classes of combinatorial K3 surfaces in (1)(-1)-form and triangulations of the sphere ${\mathbb{S}}^{2}$ such that no vertex has valency greater than $6$. In particular, a combinatorial K3 surface in (1)(-1)-form is uniquely determined by its dual intersection complex.∎
@@
Let $G$ be a triangulation of the sphere ${\mathbb{S}}^{2}$. We denote by $Y_{G}$ the unique combinatorial K3 surface in (1)(-1)-form with dual intersection complex $G$.
@@
The classification of combinatorial K3 surfaces in (1)(-1)-form through special triangulations of the sphere from Proposition 2.7 is a first instance of the combinatorial nature of combinatorial K3 surfaces. But not only those in (1)(-1)-form are determined by combinatorial objects (at least in degree $2$, even though this is likely to hold true in higher degree as well): a general combinatorial K3 surface is determined by its curve structure, a certain graph that captures the configuration of curves on the surface, see Definition 2.23. This fully justifies the terminology combinatorial K3 surface.
@@ 2.2. Type I flops
Assume that for some $i$ we have an irreducible curve $C_{i}\subset Y_{i}$ not contained in the double locus, whose preimage $C_{i}^{\nu}$ in the normalization $Y_{i}^{\nu}$ is a (1)(-1)-curve. Recall from Remark 2.2 that the double curves give rise to an anticanonical cycle on any combinatorial K3 surface. Hence, by adjunction, $C_{i}^{\nu}$ intersects the double locus in a single point $P_{i}\in D_{ij}$, which is not a triple point, and the intersection is transversal. Here $i=j$ is allowed, in which case the assumption is that in the normalization, $C_{i}^{\nu}$ meets one of the components of the preimage $D_{ii}^{1}\cup D_{ii}^{2}$, say $D_{ii}^{1}$. We then perform the following construction.
If $i\neq j$, we set $Y^{\prime\nu}_{k}:=Y^{\nu}_{k}$ for $k\neq i,j$. We have also natural identifications [^2] Let us point out that these identifications will preserve so-called interior special points, see Definition 3.2. This is however not relevant for the construction of the flop in the type I case. of the components $D_{ki},D_{kj}$ and $D^{\prime}_{ki},D^{\prime}_{kj}$ of the components of the anticanonical cycle. We define $Y^{\prime\nu}_{i}$ as the blow-down of $Y_{i}^{\nu}$ along the preimage of the (1)(-1)-curve $C_{i}$. We further define $Y^{\prime\nu}_{j}$ as the blow-up of $Y_{j}$ in the preimage [^3] This will be called the interior special point on $D_{ji}$ in the next section. of the point where $C_{i}$ meets $Y_{j}$.
@@
2.3. Combinatorial K3 surfaces of degree 22
In this article, we will mainly be concerned with combinatorial K3 surfaces of degree $2d=2$. Such surfaces have $d+1=3$ irreducible components, and the dual intersection complex is one of the two triangulations of ${\mathbb{S}}^{2}$ with three vertices; see Figure 2. As indicated in the figure, we will denote these triangulations by $P$ respectively $T$. Depending on its dual intersection complex, a combinatorial K3 surface of degree $2$ is thus said to be of class $P$ resp. $T$.
baabP{P}bbaaT{T}
@@ 2.3. Combinatorial K3 surfaces of degree 22
By Proposition 2.7, there are two combinatorial K3 surfaces of degree $2$ in (1)(-1)-form, one each of class $P$ resp. $T$. Following [HL22, Section 2.2], we call them $Y_{P}$ and $Y_{T}$. Let us recall their construction and some basic facts about their geometry.
@@
By [Laz08, Proposition 5.2, Lemma 5.14] (see also [HL22, Construction 2.25] for more details and precise references), there are weak del Pezzo surfaces ${\mathfrak{Y}}_{i}$, $i=1,\ldots,6$ such that any connected component of the normalization of any $Y\in{\operatorname{CK3}^{\scriptscriptstyle(-1)}}$ is isomorphic to one of those. More precisely, the ${\mathfrak{Y}}_{i}$ come with an anticanonical cycle ${\mathfrak{D}}_{i}$ having $i$ irreducible components and the components of the normalization of $Y$ are isomorphic to $({\mathfrak{Y}}_{i},{\mathfrak{D}}_{i})$ as log Calabi–Yau surfaces, see Remark 2.2. While the details of their construction can be found in the above references, we are mainly interested in their combinatorial description via curve structures. The latter will be defined in general in Section 2.4, but for the surfaces ${\mathfrak{Y}}_{i}$ it is straightforward: we define the set
[…]
@@
For us, only the surfaces ${\mathfrak{Y}}_{1}$, ${\mathfrak{Y}}_{2}$, and ${\mathfrak{Y}}_{4}$ will play a role. We will next describe them geometrically. To this end, we recall the notion of an $n$-blow-up from [HL22, Definition 2.24], for the reader’s convenience. The reader may also find it helpful to consult the references from the preceding example. We remark that the geometric description of $\mathfrak{Y}_{4}$ given here coincides with that in [HL22, Construction 2.25], whereas the descriptions of $\mathfrak{Y}_{1}$ and $\mathfrak{Y}_{2}$ differ, but the surfaces are the same.
Definition 2.16.
Let $(Y,D)$ be an anticanonical pair and let $p$ be a smooth point of $D$. If $n=1$, the $n$-blow-up of $Y$ in $p$ is the usual blow-up. For $n>1$, the $n$-blow-up of $Y$ in $p$ is defined inductively as the blow-up of the (n1)(n-1)-blow-up $\pi\colon Y^{\prime}\xrightarrow{\ \ }Y$ at the intersection point of the strict transform of $D$ with the exceptional set of $\pi$.
More generally, for an ordered set $(p_{1},\ldots,p_{k})$ of points lying in the smooth locus of $D$ and a $k$-tuple $n=(n_{1},\ldots,n_{k})$ of nonnegative integers, the $n$-blow-up of $Y$ in $(p_{1},\ldots,p_{k})$ is defined by applying the above construction independently at each $p_{j}$, with multiplicity $n_{j}$.
Example 2.17.
The surface $\mathfrak{Y}_{1}$ is obtained by performing the $8$-blow-up of $\mathbb{P}^{2}$ in an inflection point of a nodal cubic in the sense of Definition 2.16, and the boundary divisor $\mathfrak{D}_{1}$ is the strict transform of that cubic. This can be read off from Figure 3: After blowing down the eight curves that form the horizontal chain in the curve structure, we obtain a smooth rational surface of Picard rank one. It must therefore be isomorphic to $\mathbb{P}^{2}$. The strict transform of $\mathfrak{D}_{1}$ is a nodal cubic, and the remaining curve from the curve structure becomes a flex line of this nodal cubic.
Note that the construction is independent of the choice of a flex line. A nodal cubic has three inflection points (see e.g. [Oka07, Proposition 9 (2)]), but the group of automorphisms of ${\mathbb{P}}^{2}$ that fix the cubic acts transitively on the flex lines. This can easily be seen in coordinates: We can assume that the cubic is given by the equation
$x^{3}+y^{3}-3xyz=0$
A straightforward calculation gives that the flex points are [1:1:0][1:-1:0], $1:-\omega:0]$, and $1:-\omega^{2}:0]$ where $\omega$ denotes a primitive third root of unity. They form one orbit of the cyclic group generated by the automorphism $x:y:z]\mapsto[x:\omega y:\omega^{2}z]$.
Example 2.18.
We obtain the surface $\mathfrak{Y}_{2}$ from the Hirzebruch surface $\mathbb{F}_{2}$ together with two distinct sections $S_{1}$, $S_{2}$ of the ruling that are smooth curves of self-intersection $S_{1}^{2}=S_{2}^{2}=2$ and intersect in two distinct points. One selects a fiber of the ruling that does not meet the two points of $S_{1}\cap S_{2}$ and performs a $(3,3)$-blow up in the two points where the fiber meets $S_{1}$ and $S_{2}$. The resulting surface is $\mathfrak{Y}_{2}$, and $\mathfrak{D}_{2}$ is defined as the strict transform of the union of $S_{1}$ and $S_{2}$.
Again, this construction is independent of the choices made. A convenient model to verify this claim is the cone in ${\mathbb{P}}^{3}$ over a smooth plane conic. This cone is also the image of the map given by the linear system $|C_{0}+2f|$ where $C_{0}\subset{\mathbb{F}}_{2}$ denotes the unique (2)(-2)-curve and $f$ denotes a fiber of the ruling. Note that $S_{1},S_{2}\in|C_{0}+2f|$, i.e. they are obtained as hyperplane sections of the cone that do not meet the vertex. The automorphism group of ${\mathbb{P}}^{3}$ fixing the cone (equivalently, that of ${\mathbb{F}}_{2}$) acts transitively on ordered pairs of such hyperplanes. The stabilizer of such a pair acts transitively on the complement of the image of $S_{1}\cap S_{2}$ in the base curve of the ruling. Note that the latter is isomorphic to ${\mathbb{P}}^{1}\setminus\{0,\infty\}$ and the stabilizer contains a copy of ${\mathbb{G}}_{m}$ acting in the standard way on this.
Example 2.19.
To construct $\mathfrak{Y}_{4}$, one starts with the product surface $\mathbb{P}^{1}\times\mathbb{P}^{1}$, together with two distinct lines $A_{1}$ and $A_{2}$ from one ruling and two distinct lines $B_{1}$ and $B_{2}$ from the other ruling. Set $D_{4}:=A_{1}\cup A_{2}\cup B_{1}\cup B_{2}$. We choose two more lines, one from each ruling, distinct from the lines in $D_{4}$, and blow up the four points where they meet $D_{4}$. The resulting surface is $\mathfrak{Y}_{4}$, and $\mathfrak{D}_{4}$ is the strict transform of $D_{4}$. Note that all choices made are equivalent under the automorphism group of ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$.
1-1
@@
$1-1$
Figure 4. The augmented curve structure of the surface ${\mathfrak{Y}}_{2}$.
$1-11-1$
@@
The surfaces ${\mathfrak{Y}}_{1}$, ${\mathfrak{Y}}_{2}$, and ${\mathfrak{Y}}_{4}$ show up in the construction of the models in (1)(-1)-form, $Y_{P}$ and $Y_{T}$, as we will explain next.
@@
The surface $Y_{P}$ has three identical irreducible components $Y_{1}=Y_{2}=Y_{3}={\mathfrak{Y}}_{2}$, whose augmented curve structure is shown in Figure 4. The black vertices represent the components of the anticanonical cycle and the white vertices have self-intersection (2)(-2) unless otherwise indicated. Call $D_{ij}$, $D_{ik}$ with $\{i,j,k\}=\{1,2,3\}$ the boundary components of $Y_{i}$. We denote the singular points of the anticanonical cycle on $Y_{i}$ by $Q_{i1}$ and $Q_{i2}$. We note that there is an automorphism of $Y_{i}$ which interchanges the components $D_{ij}$ and $D_{ik}$, while leaving the points $Q_{i1}$ and $Q_{i2}$ fixed (as a set). Similarly, there is an automorphism leaving the components $D_{ij}$ and $D_{ik}$ fixed (as sets), while swapping the points $Q_{i1}$ and $Q_{i2}$. The existence of these automorphisms can, for example, be deduced from the description given in Example 2.18. In terms of the ruled surface $\mathbb{F}_{2}$ this corresponds to an automorphism interchanging $S_{1}$ and $S_{2}$, but fixing their intersection points and the ruling (as a set), or, vice versa, leaving $S_{1}$, $S_{2}$ and the ruling fixed (as sets) and interchanging the two points of intersection. In this way we obtain a group $\mu_{2}\times\mu_{2}$ of automorphisms acting on the surface $\mathfrak{Y}_{2}$.
Further, let $P_{ij}\in D_{ij}$ denote the points corresponding to the intersections of the white (1)(-1)-vertices in the curve structure with their adjacent black vertices. These points will be called interior special points, see Definition 3.2.
We can now describe how to construct the surface $Y_{P}$. We start with an ordered triple $Y_{i}$, where $i=1,2,3$, of components isomorphic to $\mathfrak{Y}_{2}$. For each component $Y_{i}$, we fix a labelling $D_{i,j}$ for $j\in\{1,2,3\}\setminus\{i\}$ of the components of the anticanonical cycle and the singular points $Q_{i,k}$ for $k=1,2$ of this cycle. The surface $Y_{P}$ is now obtained by gluing the $Y_{i}$ along certain isomorphisms of the components of the anticanonical cycles. Since these components are smooth rational curves, giving isomorphisms between them is tantamount to specifying ordered $3$-tuples of points on them that are to be identified under the isomorphism. More precisely, we choose isomorphisms $D_{12}\cong D_{21}$, $D_{23}\cong D_{32}$ and $D_{31}\cong D_{13}$ which are specified by mapping $Q_{i,k}$ to $Q_{i+1,k}$ for $i=1,2,3$ and $j=1,2$ as well as sending interior special points to interior special points. Using the symmetry groups $\mu_{2}\times\mu_{2}$ of the surfaces $Y_{i}$, which we have described above, shows that the resulting surface $Y_{P}$ is independent of any choices that have been made for the labelling of the components $D_{ij}$ and the singular points $Q_{ik}$.
@@
The surface $Y_{T}$ has irreducible components $Y_{1}=Y_{2}={\mathfrak{Y}}_{1}$ and $Y_{3}$ a nonnormal surface with normalization ${\mathfrak{Y}}_{4}$. We call $Y_{3}$ the special component and $Y_{1}$, $Y_{2}$ the nonspecial components. The curve structures of $Y_{1}=Y_{2}={\mathfrak{Y}}_{1}$ and $(Y_{3})^{\nu}={\mathfrak{Y}}_{4}$ are shown in Figure 3 and Figure 5, respectively. To obtain the surface $Y_{T}$, one glues the components as follows. First, one glues two opposite boundary components to one another, e.g. those represented by the black vertices in the bottom row of Figure 5, such that the intersection points with the remaining boundary components are matched and also the intersection points with the curves corresponding to unique white vertices adjacent to the glued boundary components are matched. This specifies the gluing uniquely. After this first gluing, the remaining boundary components, i.e. those represented by the black vertices in the top row of Figure 5, have become nodal rational curves. These are then glued to the boundary curves of the nonspecial components, again matching singular points and intersection points with neighboring curves in the curve structure.